1 * Consider the following joint distribution for the weather in two consecutive days. Let X and Y be the random variables for the weather in the first and the second days, whereas the weather is coded as 0 for sunny, 1 for cloudy, and 2 for rainy. 0 0.2 0.2 0.2 10.1 0.1 0.1 2 0 0.1 0 (a) Find the marginal probability mass functions for X and Y (b) Are the weather in two consecutive days independent? (c) Calculate the expectation and variance for X and Y (d) Calculate the covariance and correlation between X and Y. Are they correlated?
Answer #1

below is joint probability distribution of X and Y:

X
Y 0 1 2 Total
0 0.2000 0.2000 0.2000 0.6000
1 0.1000 0.1000 0.1000 0.3000
2 0.0000 0.1000 0.0000 0.1000
Total 0.3000 0.4000 0.3000 1.0000

a)

marginal distribution of X:

X P(X) XP(X) X^2P(X)
0 0.3000 0.0000 0.0000
1 0.4000 0.4000 0.4000
2 0.3000 0.6000 1.2000
total 1.0000 1.0000 1.6000
E(X) = 1.0000
E(X^2) = 1.6000
Var(X) E(X^2)-(E(X))^2 0.6000

marginal distribution of Y:

Y P(Y) YP(Y) Y^2P(Y)
0 0.6000 0.0000 0.0000
1 0.3000 0.3000 0.3000
2 0.1000 0.2000 0.4000
total 1 0.5 0.7
E(Y) = 0.5000
E(Y^2) = 0.7000
Var(Y) E(Y^2)-(E(Y))^2 0.4500

b)

here as P(X=0)*P(Y=0)=0.3*0.6 =0.18 ; which is not equal to P(X=0,Y=0)=0.20 ; hence X and Y are not independent

c)

expectation of X =E(X)=1

expectation of Y =E(Y)=0.5

Variance of X =Var(X)=0.6

Variance of Y =Var(Y)=0.45

d)

here E(XY)

=\tiny \sumxyP(x,y)=0*0*0.2+1*0*0.2+2*0*0.2+0*1*0.1+1*1*0.1+2*1*0.1+0*2*0+1*2*0.1+2*2*0=0.5

hence Covariance =Cov(X,Y)=E(XY)-E(X)E(Y)=0.0000

therefore Correlation =Cov(X,Y)/(Var(X)*Var(Y))1/2 =0

as correlation between X and Y is 0 ; threfore X and Y are uncorrelated